### v1

@@ 3.16 Masking Strategy, Model Architecture and Objective Function
A handful of works contributed to advancements in the masking strategy, the model architecture and the objective function. Half of these papers introduced a novel masking strategy, continuing with improvements that are developed on top of it [28, 178]. The other half is about initially transforming the input data, and then introducing all the modifications to apply MIM for the new input source [179, 180].
@@
Adaptive Masking (AdaMAE) [28] presented an adaptive masking strategy for MAE performed by an additional neural network that assigns greater masking probabilities to the patches containing spatio-temporal information (a.k.a. foreground). The new neural network gives a vector of probabilities from which the sampling is performed. Thus, it cannot be trained with the reconstruction loss. The solution for this problem was to use an additional loss function based on the reconstruction one, where the terms are weighted with the probability vectors given by the adaptive network.
@@
[178] presented a masking pre-training method specially designed for images with text. Random patches that contain text are masked, then passed through a CNN-based model to extract a feature representation. The feature maps are tokenized and then fed into a transformer. Following the Feature Pyramid Network (FPN) architecture [181], the resulting embedding is upsampled and combined with features maps from multiple levels. The first objective is to predict the masked words, while the second goal is to reconstruct the corrupted pixels (with the help of the predicted word tokens). A ROI-alignment unit is used to associate the feature maps with the masked regions.
@@
In order to pre-train a ViT for videos, [179] proposed to first transform the input video into the latent space of a VQ-VAE. Then, each frame is transformed into a set of tokens, and a varying high ratio of the tokens from the whole video sequence are masked. The attention units of the model are modified such that every token has access either to only the surrounding tokens in the same frame, or to a small neighboring region of tokens along all dimensions (i.e. including time). The objective is to estimate the masked tokens by minimizing their negative log-likelihood.
@@
In their work, [180] introduced a general framework to learn various computer vision tasks with transformers. They framed the inputs and output of each task (e.g. detection and segmentation) as a sequence, and used an encoder-decoder transformer with bidirectional attention masks. To capture a rich context of the task, they employed MAE pre-training by reconstructing the sequence of tokens.
@@
[182] analyzed the drawbacks in previous latent space MIM applications [19], showing how conventional reconstruction loss restricts diverse latent learning. To counter this, the authors introduced a patch discrimination objective to increase similarity between the predicted and the corresponding target latents. Additionally, [182] tackled the issue of patch correlation by changing the masking strategy. They used a high masking ratio (90%), a gap between adjacent patches, and a similarity constraint on visible and target patch sets. Further, the last contribution was an improved decoder architecture, suitable for latent representation prediction and incorporating self-attention, cross-attention layers, and visual cues from visible patches.
@@ 3.17 Masking Strategy, Target Features and Objective Function
Several papers that had an impact on the masking strategy, the target features, as well as the objective function. The contributions to the masking strategies are highly varied: integrating new target features [76, 183], improving the masking policy based on guidance [184], and even leveraging multimodal data [185]. The objective function of these works is modified as a result of the employed features.
Within the context of few-shot learning, [76] employed a masked autoencoder for reconstructing the latent embeddings rather than the original image. Additionally, each instance acts as a patch, and thus, the input consists of more images from the same class. After encoding the input, a high portion of it is masked and the decoder tries to reconstruct the masked embedding representation (given some identification variables). In is way, the backbone, i.e. the encoder, learns more discriminative features and has a better few-shot performance.
@@
[185] focused on implementing a masking pre-training strategy for both visual and textual data. The main idea is to mask one modality and reconstruct the missing data using the other input type. First, each modality is encoded with its own encoder, and further processed by its specific cross-attention encoder (which is also fed with the other modality embedding). Finally, the masked tokens are decoded using a transformer for images, or a linear classification layer for text. Besides the reconstruction objective, two other losses are employed to align the embeddings of the modalities.
@@ 3.18 Masking Strategy, Model Architecture, Downstream Task and Objective Function
A number of papers have contributed on four directions, having an extensive scope: the masking strategy, the model architecture, the downstream task and the objective function. The first paper involved integrating another learning paradigm (reinforcement learning) [186], while the rest deal with multimodal data [187, 188, 189].
[186] proposed a Token-Critic algorithm to guide the synthesis of a non-autoregressive image generation model by predicting which tokens need to be sampled or not. To train the model, the following procedure is employed: a tokenized image (through a Vector-Quantized Autoencoder) is randomly masked. Then, using the transformer-based generative model, it is reconstructed, while the critic must distinguish between the sampled and the original tokens. In this way, a completely masked tokenized image is gradually unmasked, using the Token-Critic to select which tokens to sample, eventually generating a new synthesized image.
[187] harnessed MAE to pre-train an audio-video model by integrating all prominent objectives of MIM into a two-stage framework. The first stage consists of reconstructing the original inputs, while the second stage adopts the teacher-student distillation scheme, where the latter’s objective is to reconstruct the former’s prediction. The teacher receives the full visible inputs, while the student is fed with the masked modalities. During both training stages, two different masked views of the same modality are generated, encoded, and two contrastive losses are computed: between embeddings of the same modality, as well as across modalities. Then, a joint encoder fuses the two modality embeddings, in the end being decoded with separate decoders.
@@
[189] enhanced self-supervised representation learning by leveraging audiovisual data. They proposed various pre-training architectures and objectives within a masked autoencoding framework to improve performance on audiovisual downstream classification tasks. The framework also supports multiple unimodal downstream tasks, using a single audiovisual pre-trained model.
@@ 4 Automatic Clustering
By analyzing the dendrogram, we identify several relevant clusters, which are annotated in Figure 17. The first observed category of clusters is related to the input data type: temporal data, 3D data or 3D point clouds, video, audio, and even multimodal. Other identified clusters are represented by the domain in which the MAE framework was used, or by the downstream task it was applied on: medical imaging, anomaly detection, image classification in few-shot scenarios, object detection and semantic segmentation. The clustering algorithm was also able to capture more complex concepts: training at test time, multi-view masked reconstruction, and domain or out-of-distribution adaptation. Probably one of the most notable clusters is formed by the papers that adopted the teacher-student MAE framework based on a contrastive objective. Another category consists of methods that employed diffusion models. Finally, one cluster related to the input masking strategy overlaps with our manual taxonomy.
@@
Figure 20: Sample images from MS-COCO dataset. Courtesy of [190].
@@ 5 Datasets
Various datasets have been used by different masked image modeling frameworks. Some of the most representative datasets are: CIFAR-100, ImageNet-1K, MS-COCO, UCF101, ShapeNet, CC3M, FFHQ, LAION-400M, and Visual Genome. A brief description of these datasets are provided in the following part.
@@
CIFAR-100 [193] contains low-resolution images (32$\times$32), with 600 images for each of the 100 available classes. The classes are grouped into 20 categories, called super-classes. The labels range from animals to humans and objects. CIFAR-100 is typically used to demonstrate MIM methods on downstream tasks. Some sample images from this dataset are shown in Figure 18.
@@
MS-COCO [190] contains around 330k images, in which over 1.5 million object instances are annotated across 80 object categories. The dataset includes annotations for a wide variety of tasks: object detection, segmentation, and captioning. Some sample images from this dataset are shown in Figure 20.
@@
CC3M [197] consists of approximately 3.3 million image pairs and their associated captions, describing the visual content of the images.
@@
Images ImageNet-1K [Deng et al. [2]] Pre-training & Downstream $1,281,167$ $100,000$
ImageNet-21K [Ridnik et al. [199]] Pre-training $14,197,122$
LAION-400M [Schuhmann et al. [192]] Pre-training $400,000,000$
CIFAR-100 [Krizhevsky [193]] Downstream $60,000$ $10,000$
Food-101 [Bossard et al. [200]] Downstream $75,750$ $25,250$
MS-COCO [Lin et al. [190]] Downstream $165,482$ $81,434$
ADE20K [Zhou et al. [201]] Downstream $25,574$ $2,000$
FFHQ [Karras et al. [191]] Downstream $60,000$ $10,000$
Language & Images CC3M [Sharma et al. [197]] Pre-training $,33$ 318 , 333 -
CC12M [Changpinyo et al. [202]] Pre-training $12,423,374$
MS-COCO [Lin et al. [190]] Pre-training $165,482$
Visual Genome [Krishna et al. [198]] Pre-training $108,000$
SBU Captions [Ordonez et al. [203]] Pre-training $860,000$
Video SSv2 [Goyal et al. [204]] Pre-training & Downstream $169,000$ $25,000$
Kinetics-400 [Kay et al. [195]] Pre-training & Downstream $240,000$ $20,000$
Kinetics-600 [Carreira et al. [205]] Pre-training & Downstream $390,000$ $60,000$
Kinetics-700 [Carreira et al. [206]] Pre-training & Downstream $545,317$ $105,000$
Kinetics-710 [Li et al. [207]] Pre-training $660,000$
WebVid2M [Bain et al. [208]] Pre-training $2,500,000$
UnlabeledHybrid [Wang et al. [41]] Pre-training $1,350,000$
UCF101 [Soomro et al. [194]] Downstream $9,500$ $3,500$
AVA [Li et al. [209]] Downstream $211,000$ $57,000$
3D Data ScanNet [Dai et al. [210]] Pre-training $1,513$
ShapeNet [Chang et al. [196]] Pre-training $57,448$
nuScenes [Caesar et al. [211]] Pre-training & Downstream $750$ $150$
ModelNet40 [Wu et al. [212]] Downstream $9,843$ $2,468$
ScanObjectNN [Uy et al. [213]] Downstream $11,416$ $2,882$
ShapeNetPart [Yi et al. [214]] Downstream $14,007$ $2,874$
Medical Images UKB [Spitzer et al. [215]] Pre-training $155,238$
ROCO [Pelka et al. [216]] Pre-training $81,000$
MedICaT [Subramanian et al. [217]] Pre-training $217,000$
TCIA-COVID19 [Harmon et al. [218]] Pre-training $771$
BraTS [Simpson et al. [219]] Pre-training & Downstream $351$ $191$
BTCV [Landman et al. [220]] Pre-training & Downstream $24$ $6$
APTOS [Kaggle [221]] Downstream $28,100$ $7,026$
RFMiD [Panchal et al. [222]] Downstream $2,560$ $640$
VQA-RAD [Lau et al. [223]] Downstream $3,064$ $451$
VQA-2019 [Ben Abacha et al. [224]] Downstream $12,792$ $500$
@@ 6 Performance Overview
We next provide an in-depth analysis of the performance achieved by MIM pre-training methods, when applied on leading computer vision architectures and popular benchmarks. We focus our analysis on three widely used datasets in MIM research: ImageNet-1K, MS-COCO, and Kinetics-400.
In Table 2, we present the results of different frameworks on ImageNet-1K. Most of these works obtained accuracy improvements by introducing novel masking strategies [47, 1, 15] or target features [171, 47]. Other works [49, 71] studied approaches to scale up the architectures and the number of samples used during pre-training. In contrast to these studies, some works [52] tried to limit the computational requirements of the networks, while preserving the performance of the previous works. Another significant observation that stems out from Table 2 is that integrating MIM with generative models results in substantial performance enhancements, as presented in some recent studies [128, 139]. This finding supports the widely known view that generative tasks yield robust world representations. Moreover, it demonstrates the orthogonality between MIM and generative models, suggesting complementary benefits when combined.
@@
ViT-B IN1K MAGE [Li et al. [159]] 82.5
MAGE-C [Li et al. [159]] 82.9
U-MAE [Zhang et al. [26]] 83.0
MAE [He et al. [1]] 83.6
SimMIM [Xie et al. [15]] 83.8
LocalMIM [Wang et al. [122]] 84.0
MaskFeat [Wei et al. [171]] 84.0
DMAE-B [Bai et al. [65]] 84.0
HPM [Wang et al. [47]] 84.2
GAN-MAE [Fei et al. [128]] 84.3
SemMAE [Li et al. [37]] 84.5
DiffMAE [Wei et al. [139]] 84.9
MCMAE [Gao et al. [114]] 85.0
MaskAlign [Xue et al. [129]] 85.4
Laion-50M RILS [Yang et al. [71]] 83.6
IN1K+SSv2 OmniMAE [Girdhar et al. [49]] 83.0
ViT-L IN1K U-MAE [Zhang et al. [26]] 83.2
MAGE [Li et al. [159]] 83.9
MAGE-C [Li et al. [159]] 84.3
MaskFeat [Wei et al. [171]] 85.7
HPM [Wang et al. [47]] 85.8
LocalMIM [Wang et al. [122]] 85.8
MAE [He et al. [1]] 85.9
GAN-MAE [Fei et al. [128]] 86.1
DiffMAE [Wei et al. [139]] 86.9
IN1K+SSv2 OmniMAE [Girdhar et al. [49]] 85.2
ViT-H IN1K MAE [He et al. [1]] 86.9
DiffMAE [Wei et al. [139]] 88.0
IN1K+SSv2 OmniMAE [Girdhar et al. [49]] 86.6
ViT-H448 IN1K MAE [He et al. [1]] 87.8
Swin-B IN1K GreenMIM [Huang et al. [146]] 83.8
SimMIM [Xie et al. [15]] 84.0
LocalMIM [Wang et al. [122]] 84.1
MixMAE [Liu et al. [50]] 84.6
Swin-L IN1K GreenMIM [Huang et al. [146]] 85.1
SimMIM [Xie et al. [15]] 85.4
LocalMIM [Wang et al. [122]] 85.6
SwinV2-H IN1K SimMIM [Xie et al. [15]] 85.7
SwinV2-G IN21K SimMIM [Xie et al. [15]] 90.2
ConvNeXt V2-B IN1K FCMAE [Woo et al. [112]] 84.9
ConvNeXt V2-L IN1K FCMAE [Woo et al. [112]] 85.8
ConvNeXt V2-H IN1K FCMAE [Woo et al. [112]] 86.3
In Table 3, we present a performance overview of different MIM frameworks on the MS-COCO dataset. For object detection, we report the mean Average Precision for bounding boxes (mAPbox), and for segmentation tasks, we provide the mean Average Precision for masks (mAPmask). EVA [63] stands out as the most effective method, concentrating on the expansion of training examples and network parameters. According to Table 3, its pre-training dataset integrates four distinct datasets, collectively summing over 29 million images. The closest result to EVA is obtained by SimMIM [15], when applied to a similar size architecture (SwinV2-G), but with much fewer images in the pre-training phase. As demonstrated in Table 3, the MAE strategy exhibits suboptimal performance when compared with alternative approaches applied to ViT-B and ViT-L models. Notably, among the methods outperforming MAE on ViT-B, two incorporate convolutional layers into their frameworks [96, 114]. This observation suggests that employing a hybrid architecture that combines both transformer and convolutional layers may offer significant improvements in object detection and segmentation.
Table 3: Performance on MS-COCO of different MIM pre-training schemes.
@@
ViT-Tiny IN1K MAE [He et al. [1]] 38.9 35.1
SparseMAE [Zhou et al. [115]] 47.1 42.0
ViT-B Laion-20M RILS [Yang et al. [71]] 48.5 42.6
IN1K GAN-MAE [Fei et al. [128]] 49.0 43.8
MAE [He et al. [1]] 50.3 44.9
MaskAlign [Xue et al. [129]] 52.1 45.7
MCMAE [Gao et al. [114]] 52.5 46.5
ViT-L IN1K MAE [He et al. [1]] 53.3 47.2
DiffMAE [Wei et al. [139]] 55.3 49.0
ViT-G IN21K EVA [Fang et al. [63]] 64.2 55.0
@@
Swin-T IN1K MST [Li et al. [14]] 42.7 38.8
Swin-B IN1K GreenMIM [Huang et al. [146]] 50.0 44.1
SimMIM [Xie et al. [15]] 50.4 44.4
LocalMim [Wang et al. [122]] 50.7 44.9
SwinV2-G IN22K SimMIM [Xie et al. [15]] 63.1 54.4
MIMDET-Base [Fang et al. [96]] IN1K MAE [He et al. [1]] 51.7 46.1
MIMDET-Large [Fang et al. [96]] 54.3 48.2
Results on video recognition are included in Table 4. We selected the Kinetics-400 dataset for this analysis because it was the most frequently used video dataset across the reviewed studies. A common practice among the studies listed in Table 4 is their incorporation of large-scale image datasets during the pre-training phase, alongside a video dataset. Additionally, an analysis of the performance outcomes from EVA [63] and VideoMAEv2 [41] on the ViT-G architecture reveals that the size of the pre-training dataset plays an important role in determining the final performance of the model. Larger datasets tend to provide more effective and generalized capabilities in complex video processing tasks.
Consistent with what we observed in image classification, the integration of diffusion models and MIM proves beneficial in the video domain as well. Notably, the DiffMAE model [139] achieves results that are competitive with those of EVA [63], although DiffMAE is operating with a substantially smaller model size. This finding underscores again the effectiveness of combining diffusion models with MIM techniques. However, a significant factor contributing to this performance enhancement is the utilization of a large dataset, WIT400M, during the pre-training phase. When the model is pre-trained solely with the smaller Kinetics-400 dataset, its results fall short of those achieved by VideoMAE [40]. This underscores the critical importance of employing large-scale datasets in the pre-training phase to maximize model effectiveness.
@@
ViT-B K400 VideoMAE [Tong et al. [40]] 81.5
ST-MAE [Feichtenhofer et al. [92]] 81.3
MME [Sun et al. [70]] 81.8
MGMAE [Huang et al. [52]] 81.8
OmniMAE [Girdhar et al. [49]] 80.8
K400 MVD [Wang et al. [168]] 83.4
@@
UnlabeledHybrid VideoMAEv2 [Wang et al. [41]] 81.5
ViT-L K400 VideoMAE [Tong et al. [40]] 85.2
ST-MAE [Feichtenhofer et al. [92]] 84.8
DiffMAE [Wei et al. [139]] 84.5
K400 MVD [Wang et al. [168]] 86.4
@@
UnlabeledHybrid VideoMAEv2 [Wang et al. [41]] 85.4
ViT-H K400 VideoMAE [Tong et al. [40]] 86.6
ST-MAE [Feichtenhofer et al. [92]] 85.1
K400 MVD [Wang et al. [168]] 87.3

### v2

@@ 3.16 Masking Strategy, Model Architecture and Objective Function
A handful of works contributed to advancements in the masking strategy, the model architecture and the objective function. Some of these studies leveraged their three-fold contributions to tackle unique pre-training scenarios [192, 193]. For example, [192] presented a custom solution for images that contain text. Additionally, other studies discuss about the suboptimal choices made in previous MIM pipelines [194, 195] and try to improve them. For example, [195] imposed certain properties on the masked patches to avoid the patch correlation issue.
@@
Adaptive Masking (AdaMAE) [194] employs an adaptive masking strategy for MAE performed by an additional neural network that assigns greater masking probabilities to the patches containing spatio-temporal information (a.k.a. foreground). The new neural network gives a vector of probabilities from which the sampling is performed. Thus, it cannot be trained with the reconstruction loss. The solution for this problem is to use an additional loss function based on the reconstruction one, where the terms are weighted with the probability vectors given by the adaptive network.
@@
[192] presented a masking pre-training method specially designed for images with text. Random patches that contain text are masked, then passed through a CNN-based model to extract a feature representation. The feature maps are tokenized and then fed into a transformer. Following the Feature Pyramid Network (FPN) architecture [196], the resulting embedding is upsampled and combined with feature maps from multiple levels. The first objective is to predict the masked words, while the second goal is to reconstruct the corrupted pixels (with the help of the predicted word tokens). A ROI-alignment unit is used to associate the feature maps with the masked regions.
@@
In order to pre-train a ViT for videos, [193] proposed to first transform the input video into the latent space of a VQ-VAE. Then, each frame is transformed into a set of tokens, and a varying high ratio of the tokens from the whole video sequence is masked. The attention units of the model are modified such that every token has access either to only the surrounding tokens in the same frame, or to a small neighboring region of tokens along all dimensions (i.e. including time). The objective is to estimate the masked tokens by minimizing their negative log-likelihood.
@@
In their work, [8] introduced a general framework to learn various computer vision tasks with transformers. They framed the input and output of each task (e.g. detection and segmentation) as a sequence, and used an encoder-decoder transformer with bidirectional attention masks. To capture a rich context for each task, they employed MAE pre-training by reconstructing the sequence of tokens.
@@
[195] analyzed the drawbacks in previous latent-space MIM applications [3], showing how a conventional reconstruction loss restricts diverse latent learning. To counter this, the authors introduced a patch discrimination objective to increase similarity between the predicted and the corresponding target latents. Additionally, [195] tackled the issue of patch correlation by changing the masking strategy. They used a high masking ratio (90%), a gap between adjacent patches, and a similarity constraint on visible and target patch sets. Further, the last contribution was an improved decoder architecture, suitable for latent representation prediction and incorporating self-attention, cross-attention layers, and visual cues from visible patches.
@@ 3.17 Masking Strategy, Target Features and Objective Function
Some papers had an impact on the masking strategy, the target features, as well as the objective function. The contributions to the masking strategies are highly varied: integrating new target features [197, 74], improving the masking policy based on guidance [198], and even leveraging multimodal data [199]. The objective function of these works is adapted as a result of the employed features.
Within the context of few-shot learning, [197] employed a masked autoencoder for reconstructing the latent embeddings rather than the original image. Additionally, each instance acts as a patch, and thus, the input consists of more images from the same class. After encoding the input, a high proportion of the input is masked and the decoder tries to reconstruct the masked embedding representation (given some identification variables). In this way, the backbone, i.e. the encoder, learns more discriminative features and has a better few-shot performance.
@@
[199] focused on implementing a masking pre-training strategy for both visual and textual data. The main idea is to mask one modality and reconstruct the missing data using the other input type. First, each modality is encoded with its own encoder, and further processed by its specific cross-attention encoder (which is also fed with the other modality embedding). Finally, the masked tokens are decoded using a transformer for images, or a linear classification layer for text. Besides the reconstruction objective, two more losses are employed to align the embeddings of the modalities.
@@ 3.18 Masking Strategy, Model Architecture, Downstream Task and Objective Function
Few papers have contributed in four directions, having an extensive scope: the masking strategy, the model architecture, the downstream task and the objective function. Among these, one study stands out for its integration within a reinforcement learning pipeline [200], illustrating how MIM principles can be extended beyond conventional computer vision tasks. The remaining works delve into multimodal data [80, 103, 201], proposing innovative ways of combining diverse input sources, while simultaneously refining the pre-training methodology and the architectural design.
[200] proposed a Token-Critic algorithm to guide the synthesis of a non-autoregressive image generation model by predicting which tokens need to be sampled or not. To train the model, the following procedure is employed. First, a tokenized image (through a Vector-Quantized Autoencoder) is randomly masked. Then, using the transformer-based generative model, the image is reconstructed, while the critic must distinguish between the sampled and the original tokens. In this way, a completely masked tokenized image is gradually unmasked, using the Token-Critic to select which tokens to sample, eventually generating a new synthesized image.
[80] harnessed MAE to pre-train an audio-video model by integrating all prominent objectives of MIM into a two-stage framework. The first stage consists of reconstructing the original inputs, while the second stage adopts the teacher-student distillation scheme, where the objective of the student is to reconstruct the predictions of the teacher. The teacher receives the full visible inputs, while the student is fed with the masked modalities. During both training stages, two different masked views of the same modality are generated, encoded, and two contrastive losses are computed: between embeddings of the same modality, as well as across modalities. Then, a joint encoder fuses the two modality embeddings, in the end being decoded with separate decoders. [201] also harnessed audiovisual data to enhance self-supervised representation learning. They proposed various pre-training architectures and objectives within a masked autoencoding framework to improve performance on audiovisual downstream classification tasks. The framework also supports multiple unimodal downstream tasks, using a single audiovisual pre-trained model.
@@ 4 Automatic Clustering
By analyzing the dendrogram, we identify several relevant clusters, which are annotated in Figure 20. The first observed category of clusters is related to the input data type: temporal data, 3D data or 3D point clouds, video, audio, and even multimodal. Other identified clusters are represented by the domain in which the MAE framework was used, or by the downstream task it was applied on: medical imaging, anomaly detection, image classification in few-shot scenarios, object detection, object tracking and semantic segmentation. The clustering algorithm was also able to capture more complex concepts: training at test time, multi-view masked reconstruction, and domain or out-of-distribution adaptation. Probably one of the most notable clusters is formed by the papers that adopted the teacher-student MAE framework based on a contrastive objective. Another category consists of methods that employed diffusion models. Finally, two clusters related to the input masking strategy and theoretical analysis overlap with our manual taxonomy.
@@
Figure 23: Sample images from MS COCO dataset. Courtesy of [202].
@@ 5 Datasets
Various datasets have been used by different masked image modeling frameworks. Some of the most representative datasets are: CIFAR-100, ImageNet-1K, MS COCO, UCF101, ShapeNet, CC3M, FFHQ, LAION-400M, and Visual Genome. A brief description of these datasets are provided in the following part.
@@
CIFAR-100 [205] contains low-resolution images (32$\times$32 pixels), with 600 images for each of the 100 available classes. The classes are grouped into 20 categories, called super-classes. The labels range from animals to humans and objects. CIFAR-100 is typically used to demonstrate MIM methods on downstream tasks. Some sample images from this dataset are shown in Figure 21.
@@
MS COCO [202] contains around 330k images, in which over 1.5 million object instances are annotated across 80 object categories. The dataset includes annotations for a wide variety of tasks: object detection, segmentation, and captioning. Some sample images from this dataset are shown in Figure 23.
@@
CC3M [209] consists of approximately 3.3 million images and their associated captions, describing the visual content of the images.
@@
Images ImageNet-1K [10] Pre-training & Downstream $1,281,167$ $100,000$
ImageNet-21K [211] Pre-training $14,197,122$
LAION-400M [204] Pre-training $400,000,000$
CIFAR-100 [205] Downstream $60,000$ $10,000$
Food-101 [212] Downstream $75,750$ $25,250$
MS COCO [202] Downstream $165,482$ $81,434$
ADE20K [213] Downstream $25,574$ $2,000$
FFHQ [203] Downstream $60,000$ $10,000$
Language & Images CC3M [209] Pre-training $,33$ 318 , 333 -
CC12M [214] Pre-training $12,423,374$
MS COCO [202] Pre-training $165,482$
Visual Genome [210] Pre-training $108,000$
SBU Captions [215] Pre-training $860,000$
Video SSv2 [216] Pre-training & Downstream $169,000$ $25,000$
Kinetics-400 [207] Pre-training & Downstream $240,000$ $20,000$
Kinetics-600 [217] Pre-training & Downstream $390,000$ $60,000$
Kinetics-700 [218] Pre-training & Downstream $545,317$ $105,000$
Kinetics-710 [219] Pre-training $660,000$
WebVid2M [220] Pre-training $2,500,000$
UnlabeledHybrid [57] Pre-training $1,350,000$
UCF101 [206] Downstream $9,500$ $3,500$
AVA [221] Downstream $211,000$ $57,000$
3D Data ScanNet [222] Pre-training $1,513$
ShapeNet [208] Pre-training $57,448$
nuScenes [223] Pre-training & Downstream $750$ $150$
ModelNet40 [224] Downstream $9,843$ $2,468$
ScanObjectNN [225] Downstream $11,416$ $2,882$
ShapeNetPart [226] Downstream $14,007$ $2,874$
Medical Images UKB [227] Pre-training $155,238$
ROCO [228] Pre-training $81,000$
MedICaT [229] Pre-training $217,000$
TCIA-COVID19 [230] Pre-training $771$
BraTS [231] Pre-training & Downstream $351$ $191$
BTCV [232] Pre-training & Downstream $24$ $6$
APTOS [233] Downstream $28,100$ $7,026$
RFMiD [234] Downstream $2,560$ $640$
VQA-RAD [235] Downstream $3,064$ $451$
VQA-2019 [236] Downstream $12,792$ $500$
@@ 6 Performance Overview
We next provide an in-depth analysis of the performance achieved by MIM pre-training methods, when applied on leading computer vision architectures and popular benchmarks. We focus our analysis on three widely used datasets in MIM research: ImageNet-1K, MS COCO, and Kinetics-400.
In Table 2, we present the results of different frameworks on ImageNet-1K. Most of these works obtained accuracy improvements by introducing novel masking strategies [47, 9, 2] or target features [73, 47]. In general, the new masking strategies are based on the intuition that randomly masking patches, as proposed in MAE [9] and SimMIM [2], is suboptimal. Thus, recent methods [47, 42, 51] aim to identify semantically important patches, usually implementing masking strategies in an easy-to-hard pipeline, starting by masking the least important patches and progressively targeting the more semantically significant ones. Other works [38, 82] studied approaches to scale up the architectures and the number of samples used during pre-training. In contrast to these studies, some works [58] tried to limit the computational requirements of the networks, while preserving the performance of the previous works. Another significant observation that stems out from Table 2 is that integrating MIM with generative models results in substantial performance enhancements, as presented in some recent studies [142, 156]. This finding supports the widely known view that generative tasks yield robust world representations. Moreover, it demonstrates the orthogonality between MIM and generative models, suggesting complementary benefits when combined.
@@
ViT-B IN1K MAGE [174] 82.5
MAGE-C [174] 82.9
U-MAE [34] 83.0
MAE [9] 83.6
SimMIM [2] 83.8
LocalMIM [138] 84.0
MaskFeat [73] 84.0
DMAE-B [78] 84.0
HPM [47] 84.2
GAN-MAE [142] 84.3
SemMAE [42] 84.5
DiffMAE [156] 84.9
MCMAE [128] 85.0
CMAE [4] 85.3
MaskAlign [76] 85.4
Laion-50M RILS [82] 83.6
IN1K+SSv2 OmniMAE [38] 83.0
ViT-L IN1K U-MAE [34] 83.2
MAGE [174] 83.9
MAGE-C [174] 84.3
MaskFeat [73] 85.7
HPM [47] 85.8
LocalMIM [138] 85.8
MAE [9] 85.9
GAN-MAE [142] 86.1
DiffMAE [156] 86.9
IN1K+SSv2 OmniMAE [38] 85.2
ViT-H IN1K MAE [9] 86.9
DiffMAE [156] 88.0
IN1K+SSv2 OmniMAE [38] 86.6
ViT-H448 IN1K MAE [9] 87.8
Swin-B IN1K GreenMIM [162] 83.8
SimMIM [2] 84.0
LocalMIM [138] 84.1
MixMAE [39] 84.6
Swin-L IN1K GreenMIM [162] 85.1
SimMIM [2] 85.4
LocalMIM [138] 85.6
SwinV2-H IN1K SimMIM [2] 85.7
SwinV2-G IN21K SimMIM [2] 90.2
ConvNeXt V2-B IN1K FCMAE [127] 84.9
ConvNeXt V2-L IN1K FCMAE [127] 85.8
ConvNeXt V2-H IN1K FCMAE [127] 86.3
In Table 3, we present a performance overview of different MIM frameworks on the MS COCO dataset. For object detection, we report the mean Average Precision for bounding boxes (mAPbox), and for segmentation tasks, we provide the mean Average Precision for masks (mAPmask). EVA [75] stands out as the most effective method, concentrating on the expansion of training examples and network parameters. According to Table 3, its pre-training dataset integrates four distinct datasets, collectively summing over 29 million images. The closest result to EVA is obtained by SimMIM [2], when applied to a similar size architecture (SwinV2-G), but with much fewer images in the pre-training phase. As demonstrated in Table 3, the MAE strategy exhibits suboptimal performance when compared with alternative approaches applied to ViT-B and ViT-L models. Notably, among the methods outperforming MAE on ViT-B, two incorporate convolutional layers into their frameworks [96, 128]. This observation suggests that employing a hybrid architecture that combines both transformer and convolutional layers may offer significant improvements in object detection and segmentation.
In terms of training objectives, based on the results reported in Tables 2 and 3, the reconstruction loss remains the primary driver for high performance in MIM methods across both ImageNet-1K and MS COCO. At the same time, the contrastive loss often serves as a complementary component. For example, one of the top-performing methods, CMAE [4], leverages both reconstruction and contrastive losses.
Table 3: Performance on MS COCO of different MIM pre-training schemes.
@@
ViT-Tiny IN1K MAE [9] 38.9 35.1
SparseMAE [124] 47.1 42.0
ViT-B Laion-20M RILS [82] 48.5 42.6
IN1K GAN-MAE [142] 49.0 43.8
MAE [9] 50.3 44.9
MaskAlign [76] 52.1 45.7
MCMAE [128] 52.5 46.5
CMAE [4] 52.9 47.0
ViT-L IN1K MAE [9] 53.3 47.2
DiffMAE [156] 55.3 49.0
ViT-G IN21K EVA [75] 64.2 55.0
@@
Swin-T IN1K MST [1] 42.7 38.8
Swin-B IN1K GreenMIM [162] 50.0 44.1
SimMIM [2] 50.4 44.4
LocalMim [138] 50.7 44.9
SwinV2-G IN22K SimMIM [2] 63.1 54.4
MIMDET-Base [96] IN1K MAE [9] 51.7 46.1
MIMDET-Large [96] 54.3 48.2
Results on video recognition are included in Table 4. We select the Kinetics-400 dataset for this analysis because it is the most frequently used video dataset across the reviewed studies. A common practice among the studies listed in Table 4 is their incorporation of large-scale image datasets during the pre-training phase, alongside a video dataset. Additionally, an analysis of the performance outcomes from EVA [75] and VideoMAEv2 [57] on the ViT-G architecture reveals that the size of the pre-training dataset plays an important role in determining the final performance of the model. Larger datasets tend to provide more effective and generalized capabilities in complex video processing tasks.
Consistent with what we observed in image classification, the integration of diffusion models and MIM proves beneficial in the video domain as well. Notably, the DiffMAE model [156] achieves results that are competitive with those of EVA [75], although DiffMAE is operating with a substantially smaller model size. This finding underscores again the effectiveness of combining generative diffusion models with MIM techniques. However, a significant factor contributing to this performance enhancement is the utilization of a large dataset, WIT400M, during the pre-training phase. When the model is pre-trained solely with the smaller Kinetics-400 dataset, its results fall short of those achieved by VideoMAE [35]. This underscores the critical importance of employing large-scale datasets in the pre-training phase to maximize model effectiveness.
@@
ViT-B K400 VideoMAE [35] 81.5
ST-MAE [102] 81.3
MME [70] 81.8
MGMAE [58] 81.8
OmniMAE [38] 80.8
K400 MVD [183] 83.4
@@
UnlabeledHybrid VideoMAEv2 [57] 81.5
ViT-L K400 VideoMAE [35] 85.2
ST-MAE [102] 84.8
DiffMAE [156] 84.5
K400 MVD [183] 86.4
@@
UnlabeledHybrid VideoMAEv2 [57] 85.4
ViT-H K400 VideoMAE [35] 86.6
ST-MAE [102] 85.1
K400 MVD [183] 87.3